A speeding car is travelling at a constant speed of 25m/s when it passes a stopped Car. If the police are accelerates at 6m/s squared how far will the police car have travelled when it catches the speeding car
1
Expert's answer
2019-09-18T09:46:52-0400

The equation of motion of a speeding car


"x_{1}=vt=25t"

The equation of motion of a police car


"x_2=\frac{at^2}{2}=\frac{6t^2}{2}=3t^2"

When a police car catches the speeding car


"x_1=x_2"

So


"25t=3t^2"

"t=\frac{25}{3}\:\rm s=8.3\: s"